OpenStudy (anonymous):

Two hockey pucks, labeled A and B, are initially at rest on a smooth ice surface and are separated by a distance of 18.0 . Simultaneously, each puck is given a quick push, and they begin to slide directly toward each other. Puck A moves with a speed of 1.30 , and puck B moves with a speed of 1.70 . What is the distance covered by puck A by the time the two pucks collide?

OpenStudy (anonymous):

Let puck B be placed at x=0 and A at x=18m At any instant 't', X co-ordinate of B is Xb=0+1.70t, and that of A is Xa=18+1.30t. Now, let the two pucks collide at 't', then at 't', Xa=Xb 1.70t=18+1.30t solving for 't' gives t=45 sec. Distence travelled be A in 45 sec. = 1.30*45 m = 58.5 m